REANALYSIS OF THE VISIBLE AND ULTRAVIOLET SPECTRA OF AgO

\underline{L.C. O’BRIEN}, S.J. WALL, G.L. HENRY, Department of Chemistry, Southern Illinois University at Edwardsville, Edwardsville, IL 62026-1652.

The C^{2}Π − X^{2}Π and D^{2}Π − X^{2}Π electronic transitions of AgO, recorded by Uhler (1953), have been reanalyzed. The ground state constants were held to those determined by our recent analysis of the A^{2}Σ − X^{2}Π near-IR transition. Improved excited state constants will be presented.